### v3

# VidCRAFT3: Camera, Object, and Lighting Control for Image-to-Video Generation

Abstract
Recent image-to-video generation methods have demonstrated success in enabling control over one or two visual elements, such as camera motion or object motion. However, these methods are unable to offer control over multiple visual elements due to limitations in data and network efficacy. In this paper, we introduce VidCRAFT3, a novel framework for precise image-to-video generation that enables control over camera motion, object motion, and lighting direction simultaneously. VidCRAFT3 integrates three core components: Image2Cloud generates 3D point cloud from a reference image; ObjMotionNet encodes sparse object trajectories using multi-scale optical flow features; and Spatial Triple-Attention Transformer incorporates lighting direction embeddings via parallel cross-attention modules. Additionally, we introduce the VideoLightingDirection dataset, providing synthetic yet realistic video clips with accurate per-frame lighting direction annotations, effectively mitigating the lack of annotated real-world datasets. We further adopt a three-stage training strategy, ensuring robust learning even without joint multi-element annotations. Extensive experiments show that VidCRAFT3 produces high-quality video content, outperforming state-of-the-art methods in control granularity and visual coherence. Code and data will be publicly available.
Figure 1: VidCRAFT3 is the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction. It offers user-friendly control over camera motion (a trajectory in blue), object motion (sparse trajectories in red), and lighting direction. VidCRAFT3 can take any combination of supported control signals and deliver fine-grained and faithful generation results.
1 Introduction
Image-to-video (I2V) generation is a powerful technique that brings still images to life, enabling widespread applications in the fields of content creation, advertising, and art animation. Typically, I2V generation aims to animate a reference image according to user-defined control signals, such as text, object motion, and camera motion, while maintaining the quality and natural dynamics of the generated video. Driven by advances in diffusion-based generative models and extensive web-scale data [2], recent methods have significantly enhanced the ability to generate temporally coherent and visually compelling videos from limited inputs such as a single image or sparse annotations [15, 6, 60, 14, 57, 17, 28, 13]. However, achieving precise and simultaneous control over multiple video attributes, such as camera motion, object motion, and lighting direction, remains a critical yet challenging objective.
Existing approaches typically address these control attributes independently or in a partially integrated manner. Methods focusing exclusively on camera control, such as CameraCtrl [17] and ViewCrafter, do not adequately handle detailed object motion or dynamic lighting variations. Similarly, object-motion-centric frameworks, including DragAnything [58], lack integrated capabilities for precise camera and lighting control. For lighting control, techniques like DILightNet [66] and NeuLighting [27] provide lighting control but often lack generalizability due to reliance on specific categories (e.g., human faces) and object-specific annotations. Other methods [27, 68, 22, 68, 22] also primarily target narrow domains, such as human faces, further restricting their applicability to open-domain scenarios.
Simultaneous and fine-grained control of camera motion, object motion, and lighting direction introduces multiple significant challenges: (1) Accurate camera motion control from single-image inputs requires robust extraction of sufficient 3D geometric information. (2) Realistic and detailed object motion control demands effective representation and refinement of sparse motion trajectories without compromising visual fidelity. (3) Dynamic lighting control necessitates integrating illumination adjustments coherently with both camera and object motion to ensure temporal consistency and visual realism.
To overcome these challenges, we propose VidCRAFT3, which integrates three core components. First, the Image2Cloud module leverages DUSt3R [54] to reconstruct a 3D point cloud from a single reference image, enabling precise camera motion control by rendering the point cloud along user-defined camera motion trajectory. Second, ObjMotionNet encodes sparse object motion trajectories by extracting multi-scale motion features from Gaussian-smoothed optical flow maps to guide realistic object motion. Third, Spatial Triple-Attention Transformer integrates lighting direction control by integrating lighting embedding with image-text embeddings through parallel cross-attention layers, ensuring consistent illumination effects. The capabilities of these modules are shown in Fig. 1, showcasing VidCRAFT3’s control over camera motion, object motion, and lighting direction. To address data scarcity, we introduce the VideoLightingDirection (VLD) dataset, providing synthetic yet highly realistic video sequences with accurate per-frame lighting annotations. Additionally, we develop a three-stage training strategy that enables effective learning without joint multi-element annotations.
The main contributions of this paper are: (1) VidCRAFT3 is the first I2V framework to achieve simultaneous control over camera motion, object motion, and lighting direction through a disentangled architecture combining Image2Cloud, ObjMotionNet, and Spatial Triple-Attention Transformer. (2) VLD dataset provides synthetic videos accompanied by accurate per-frame lighting direction annotations, effectively addressing the scarcity of annotated real-world datasets. (3) A three-stage training strategy enables robust multi-element control without requiring real-world videos annotated with camera, object, and lighting signals. (4) Extensive experiments demonstrate that VidCRAFT3 achieves state-of-the-art performance surpassing existing methods in terms of control precision, visual quality, and generalization capabilities.
2 Related Work
2.1 Image-to-video Generation
Image-to-video (I2V) generation [15, 6, 7, 60, 14, 16] aims to animate static images into dynamic videos while preserving visual content and introducing realistic motion. Recent advances in diffusion models [48, 20] have revolutionized video generation by extending pre-trained Text-to-Image (T2I) models like AnimateDiff [15] to incorporate temporal dimensions for motion generation. These methods integrate the input image as a condition, either through CLIP-based [42] image embeddings or by concatenating the image with noisy latent. For example, VideoCrafter1 [6], DynamiCrafter [60], and I2V-Adapter [14] use dual cross-attention layers to fuse image embeddings with noisy frames, ensuring spatial-aligned guidance. Similarly, Stable Video Diffusion (SVD) [4] replaces text embeddings with CLIP image embeddings, maintaining semantic consistency in an image-only manner. Another line of work, exemplified by SEINE [8], DynamiCrafter [60] and PixelDance [67], expands the input channels of diffusion models to concatenate the static image with noisy latents, effectively injecting image information into the model. However, these methods preserve input image fidelity while generating dynamic videos but often struggle with fine-grained details due to reliance on global conditions.
2.2 Motion-controlled Video Generation
Motion-controlled video generation [57, 17, 64] focuses on creating high-fidelity videos with user-defined motion dynamics. Existing methods can be categorized into camera motion control, object motion control, and joint motion control. For camera motion control [17, 1, 26, 21, 56], CamI2V [72], CameraCtrl [17], and CamCo [61] encode camera parameters into Plücker embeddings for more discriminative motion representation. ViewCrafter [65] and I2VControl-Camera [11] use 3D representations of reference images to render point clouds and generate partial frames. For object motion control [32, 46, 55, 23, 30, 71, 74, 62, 37, 19, 12, 34, 59], existing methods employ diverse strategies to guide object movements in video synthesis. DragNUWA [64], Image Conductor [28], DragAnything [58], and MotionBridge [50] utilize sparse optical flow to control object motion, while MOFA-Video [36] and Motion-I2V [45] predict dense motion fields to control object motion. Alternatively, Boximator [52] and Direct-a-Video [63] rely on bounding box trajectories to guide object motion. LeviTor [51] combines depth and K-means clustering for precise 3D trajectory control. Current research on Joint Motion Control [63, 57, 13, 9], aiming to simultaneously control camera and object motions, remains limited. MotionCtrl [57] designs specialized modules for precise control, while Motion Prompting [13] uses motion tracks for flexible guidance. Perception-as-Control [9] further enhances control with 3D-aware motion representations.
Unlike existing methods, we propose VidCRAFT3, the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction. By combining 3D point cloud rendering, trajectory learning, and Spatial Triple-Attention Transformer, our approach effectively decouples these elements, ensuring temporal consistency and enhanced realism in complex scenes.
2.3 Lighting-controlled Visual Generation
Lighting-controlled visual generation aims to manipulate illumination while preserving scene geometry and materials. Previous methods primarily focus on portrait lighting [49, 73, 43, 38, 35, 25, 44, 47, 18, 40, 70], laying the foundation for effective and accurate illumination modeling. Recent advances in diffusion models have significantly improved the quality and flexibility of lighting control. Methods like DiLightNet [66] and GenLit [3] achieve fine-grained and realistic relighting through radiance hints and SVD, respectively. Facial relighting methods, including DifFRelight [18] and DiFaReli [40], produce high-quality portrait images. Frameworks like NeuLighting [27] focus on outdoor scenes using unconstrained photo collections, while GSR [39] combines diffusion models with neural radiance fields for realistic 3D-aware relighting. IC-Light [69] proposes imposing consistent light transport during training. Extending lighting control to video [70] introduces challenges such as temporal consistency and dynamic lighting effects. Recent techniques leverage 3D-aware generative models for temporally consistent relighting, as seen in EdgeRelight360 [29] and ReliTalk [41]. Neural rendering approaches [5, 68] use datasets like dynamic OLAT for high-quality portrait video relighting, while reflectance field-based methods [22] infer lighting from exemplars. Despite these advancements, most prior works focus on portraits and HDR lighting conditions. In contrast, VidCRAFT3 targets general scenes, enabling interactive adjustment of lighting directions.
3 Method
3.1 Overview
Figure 2: Architecture of VidCRAFT3 for controlled image-to-video generation. The model builds on Video Diffusion Model (VDM) and consists of three main components: the Image2Cloud reconstructs 3D point cloud from a single reference image and generates point cloud renderings along user-defined camera motion trajectory; the ObjMotionNet injects object dynamics into the UNet by encoding sparse trajectories into multi-scale motion features; the Spatial Triple-Attention Transformer integrates image, text, and lighting information via parallel cross-attention modules. The model enables I2V generation conditioned on arbitrary combination of camera motion, object motion, and lighting direction.
We present VidCRAFT3, the first framework for high-fidelity Image-to-Video (I2V) generation with disentangled control over camera motion, object motion, and lighting direction. Existing methods struggle to simultaneously manage these elements due to entangled representations and insufficient annotated training data. To address this challenge, as shown in Fig. 2, our framework integrates three core components: (1) Image2Cloud, which reconstructs a 3D point cloud from a single input image and generates point cloud renderings along user-defined camera motion trajectory for precise camera control; (2) ObjMotionNet, which encodes sparse object trajectories into multi-scale motion features for realistic object dynamics; and (3) the Spatial Triple-Attention Transformer, which integrates lighting direction, image, and text features via parallel cross-attention layers. To overcome the scarcity of real-world videos with multi-element annotations, we construct three specialized datasets and adopt a three-stage training strategy to progressively optimize the model.
Our model is built upon the open-source I2V model DynamiCrafter, which consists of a VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ for image compression, a UNet with spatial and temporal layers, and CLIP text and image encoders. We employ a dual-stream injection approach where the reference image is encoded by CLIP image encoder and injected into UNet through image cross-attention. Simultaneously, the first frame of the point cloud rendering is replaced with the reference image, which is then encoded by the VAE encoder and concatenated with noise before being fed into the UNet. This design ensures seamless integration of spatial and temporal information for high-quality video generation.
3.2 Model Architecture
Camera Motion Control via Point Cloud Rendering. To achieve precise camera motion control in I2V, VidCRAFT3 leverages the Image2Cloud module, which reconstructs a high-quality 3D point cloud of the scene from a single reference image. Specifically, we employ DUSt3R, an unconstrained stereo 3D reconstruction model, to generate a 3D point cloud. Given a reference image $I_{\text{ref}}$, DUSt3R performs monocular or binocular reconstruction via point regression, followed by global alignment to ensure multi-view consistency: $\mathcal{P}=\text{DUSt3R}(I_{\text{ref}})$. The reconstructed point cloud provides explicit 3D geometry, enabling accurate rendering of the scene from arbitrary camera motion trajectory $E=\{E_{1},E_{2},\ldots,E_{F}\}$, where $F$ is the length of trajectory. Given a user-defined trajectory $E$, the point cloud rendering at frame $t$ is computed as $R_{t}=\pi(\mathcal{P},E_{t})$, where $\pi(\cdot)$ is the differentiable rendering function. However, due to the limitations of point cloud representation and the sparse 3D cues from a single image, the point cloud renderings may exhibit artifacts such as missing regions, occlusions, and geometric distortions. To address this, VidCRAFT3 integrates point cloud renderings as a input of VDM, which refines the coarse renderings to generate high-quality and temporally consistent video frames. This combination of explicit 3D geometry and generative refinement ensures both accurate camera control and realistic video synthesis.
Object Motion Control through Trajectory Learning. Object motion in VidCRAFT3 is controlled through sparse spatial trajectories. For up to $N$ objects in a $F$-frames video, each trajectory is defined as a sequence of 2D pixel coordinates $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\mid n\in[1,N],\,f\in[1,F]\}$, where $\mathbf{s}_{n}^{f}$ denotes the position of the $n$-th object in frame $f$. To model motion dynamics, we compute inter-frame optical flow vectors between consecutive frames. For each trajectory point $\mathbf{s}_{n}^{f}$, the displacement vector $\mathbf{v}_{n}^{f}$ is calculated as $\mathbf{v}_{n}^{f}=\mathbf{s}_{n}^{f+1}-\mathbf{s}_{n}^{f}=(x_{n}^{f+1}-x_{n}^{f},\,y_{n}^{f+1}-y_{n}^{f}).$ These sparse motion vectors are then projected into a per-frame optical flow map $\mathcal{V}^{f}\in\mathbb{R}^{H\times W\times 2}$, where $H$ and $W$ denote spatial dimensions. The mapping is formalized as
$\mathcal{V}^{f}(x,y)=\begin{cases}\mathbf{v}_{n}^{f}&\text{if }(x,y)=(x_{n}^{f},y_{n}^{f})\text{ for any }n,\\ (0,0)&\text{otherwise},\end{cases}$ (1)
with the first frame’s flow initialized as $\mathcal{V}^{1}(x,y)=(0,0),\forall(x,y)$. The full spatiotemporal flow tensor $\mathcal{V}\in\mathbb{R}^{F\times H\times W\times 2}$ is subsequently processed through Gaussian smoothing to obtain a dense motion representation $\tilde{\mathcal{V}}$. The ObjMotionNet is a neural network composed of multiple convolutional layers and downsampling operations, designed to extract multi-scale motion features from $\tilde{\mathcal{V}}$. Inspired by T2I-Adapter [33], ObjMotionNet injects multi-scale motion features exclusively into the UNet Encoder. This balances precise motion control with video quality, as the encoder handles structure while the decoder refines details, ensuring accurate guidance without compromising output quality.
Figure 3: Construction pipeline of the Object Motion Control Dataset.
Lighting Direction Control with Spatial Triple-Attention Transformer. The lighting direction is represented as a per-frame 3D vector $L=(l_{x},l_{y},l_{z})$, describing the orientation of light source in Cartesian coordinates. To effectively encode this directional information into a high-dimensional feature space, we employ Spherical Harmonic (SH) Encoding. SH encoding captures the angular characteristics of the lighting using basis functions up to degree 4, resulting in 16 coefficients. The resulting SH-encoded vector $L_{\text{SH}}\in\mathbb{R}^{16}$ is projected into the feature space of the UNet using a multi-layer perceptron (MLP). $\mathbf{E}_{\text{light}}=\text{MLP}(L_{\text{SH}})$, where $\mathbf{E}_{\text{light}}$ is a lighting embedding aligned with the dimensionality of text embedding.
To incorporate the lighting embedding into the UNet, we propose the Spatial Triple-Attention Transformer, which integrates three parallel attention modules: image cross-attention, text cross-attention, and lighting cross-attention. The lighting cross-attention module integrates the encoded lighting embedding $\mathbf{E}_{\text{light}}$ into the UNet. This attention mechanism modulates the spatial features based on the input lighting direction. The operation is defined as
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
where $Q$ (Query) comes from the self-attention output of the UNet, $K$ (Key) and $V$ (Value) are derived from $\mathbf{E}_{\text{light}}$. The outputs of three cross-attention are summed to produce a fused feature representation $\mathbf{O}=\mathbf{O}_{\text{image}}+\mathbf{O}_{\text{text}}+\mathbf{O}_{\text{light}}$ where $\mathbf{O}_{\text{image}}$, $\mathbf{O}_{\text{text}}$, and $\mathbf{O}_{\text{light}}$ are the outputs of the image, text, and lighting cross-attention modules, respectively. This novel mechanism ensures the generated videos maintain consistency across lighting, text, and image conditions.
Figure 4: Construction pipeline of the VLD Dataset
3.3 Dataset Construction
Due to the lack of datasets with annotations for camera motion trajectory, object motion trajectories, and lighting directions, we construct three specialized datasets. All datasets consist of 25-frame video clips with a spatial resolution of $320\times 512$ pixels.
Camera Motion Control Dataset. We constructe this dataset from RealEstate10K [75], creating 62,000 clips with smooth camera transitions. Using DUSt3R, we generated globally aligned point clouds and rendered the first frame’s point cloud along the ground-truth camera trajectory. Since RealEstate10K lacks captions, we uniformly sampled 4 frames per clip and generated descriptions using Qwen2-VL-7B-Instruct [53]. Despite its focus on indoor scenes, the dataset provides diverse camera trajectories for fine-grained motion control.
Object Motion Control Dataset. We construct this dataset from WebVid-10M [2], consisting of 60,000 video clips. The dataset construction pipeline, illustrated in Fig. 3, consists of the following five steps: (1) Clip Filtering: Clips with abrupt scene changes are removed using PySceneDetect, and sequences with temporal intervals (1–16 frames) are sampled. To retain only clips with significant motion, optical flow is computed via MemFlow [10], and the bottom 25% of clips with low motion scores are filtered out. (2) Dense Trajectory Generation: For each filtered clip, CoTrackerV3 [24] generates dense trajectories on a $16\times 16$ grid, capturing pixel-level motion dynamics across frames. (3) Trajectory Filtering by Length: Trajectories shorter than the average length are discarded to ensure meaningful object motion. (4) Sparse Trajectory Sampling: From the filtered dense trajectories, 1–8 sparse trajectories are sampled with probability proportional to their length. (5) Optical Flow Smoothing: Optical flow between adjacent frames is computed to encode motion direction and intensity. Finally, a Gaussian filter is applied to smooth the sparse trajectory matrix, ensuring stable training.
Lighting Direction Control Dataset. Since it is challenging to control a camera to follow the same trajectory under varying lighting conditions in real-world scenarios, collecting such datasets is both extremely difficult and costly. To address this, we introduce the VideoLightingDirection (VLD) Dataset, a novel synthetic dataset designed for modeling complex light-object interactions. The VLD Dataset comprises 57,600 synthetic videos rendered using Blender, featuring 3,600 realistic scenes, each with 16 videos captured under distinct lighting conditions. As illustrated in Fig. 4, the dataset construction pipeline consists of the following four steps: (1) Scene Creation: We designed two scene types: Haven scenes and BOP scenes. Haven scenes places 3D models from Poly Haven at the center of HDR environments, while BOP scenes randomly places BOP models in a six-plane textured room to simulate indirect lighting. Models and environments can be randomly combined for diversity. See supplementary materials for BOP scenes examples. (2) Camera Trajectory Sampling: We sample the starting point of the camera trajectory on a spherical region (radius: 0.7–1.3 meters) centered around the 3D models. A smooth trajectory is then randomly generated around the model, ensuring that the camera always facing the center of the model. (3) Lighting Direction Sampling: To enhance the lighting effect, we reduce the HDR environment intensity by 40%. We uniformly sample 16 points on a hemisphere centered on the model, with the base surface normal aligned with the camera’s viewing direction. Each sampled point serves as the position of a 2kW spotlight (radius: 1), oriented toward the model’s center. The light source position vectors are mapped to all camera coordinate systems and normalized to compute the lighting direction for each view. (4) Rendering and Annotation: Using Blender Cycles, we render each scene under 16 different lighting conditions while maintaining a consistent camera trajectory. Each frame is annotated with the corresponding camera pose and lighting direction, providing precise control signals for training.
3.4 Training Strategy
To enable simultaneous control over camera motion, object motion, and lighting direction, we adopt a three-stage progressive training strategy, designed to address the absence of datasets annotated with all three control signals.
Stage 1: Camera Motion Control Training. The training begins by initializing the model with DynamiCrafter pre-trained weights. The model is then fine-tuned on camera motion control dataset for 40,000 iterations, optimizing the entire UNet to align VDM with camera motion. This stage establishes robust 3D scene understanding by integrating point cloud renderings, ensuring precise global camera movements while maintaining temporal consistency.
Stage 2: Dense Object Trajectories and Lighting Mixed Fine-tuning. We combine the Object Motion Control Dataset and VLD Datasets to create a comprehensive dataset annotated with camera motion, object motion, and lighting direction. This hybrid dataset enhances the ability of model to learn joint control over these three conditions. Dense object trajectories are incorporated to provide rich motion details, accelerating model convergence. To retain the camera control capabilities from stage 1, the temporal layers of the UNet remain frozen, while the spatial layers and newly introduced components, including ObjMotionNet, lighting cross-attention, and the MLP for lighting direction projection, are optimized for 20,000 iterations. This stage enables the model to simultaneously control all three conditions while ensuring global camera alignment.
Stage 3: Sparse Object Trajectories and Lighting Mixed Fine-tuning. The same hybrid dataset from Stage 2 is reused, but dense trajectories are replaced with sparse trajectories to simulate real-world user interactions. Fine-tuning follows the same parameters as Stage 2, with 20,000 iterations. This stage forces the model to infer complex motion patterns from limited trajectory data, and the progressive shift from dense to sparse supervision enhances its ability to generalize to practical scenarios. By progressively learning camera motion, object motion, and lighting direction controls across stages while strategically freezing layers to retain prior knowledge, the training strategy ensures that VidCRAFT3 achieves fine-grained, synergistic control over all three elements, even in the absence of fully annotated multi-task training data.
4 Experiments
Figure 5: Qualitative comparisons with SOTA methods on RealEstate10K.
Figure 6: Qualitative comparisons with SOTA methods on WebVid-10M.
4.1 Experimental Setup
Implementation Details. Our model builds upon DynamiCrafter, initialized with its pre-trained weights. During training and inference, video clips are processed at $320\times 512$ resolution with 25 frames. To optimize the model, we employ the Adam optimizer with a learning rate of $1\times 10^{-5}$ and a batch size of 96. The training is conducted on 8 NVIDIA H100 GPUs. For inference, we adopt the DDIM sampler and classifier-free guidance to enhance generation quality and multimodal consistency. During inference, the model averagely uses 20 GB of GPU memory and takes 42 seconds.
[…]
CameraCtrl 97.99 96.11 29.41 4.19
MotionCtrl 103.82 188.93 30.18 4.23
CamI2V 98.54 85.03 30.37 4.24
Ours 75.62 49.77 32.32 4.07
Table 1: Quantitative comparison with SOTA methods on RealEstate10K.
Evaluation Datasets. We evaluate our model on three domain-specific datasets and a generalized test set. For camera motion control, we sample 1,000 samples from the RealEstate10K test set. For object motion control, we sample 1,000 samples from the WebVid-10M test set. For lighting direction control, we sample 1,000 samples from the VLD dataset, covering a wide range of lighting directions. To provide a broader evaluation, we create a generalized test set consisting of 100 videos sourced from copyright-free websites like Pixabay and Pexels, as well as videos generated by T2V models. This dataset spans categories such as human activities, animals, vehicles, indoor scenes, artworks, natural landscapes, and AI-generated images.
Evaluation Metrics. We evaluate VidCRAFT3 across three dimensions: (1) Video Quality: FID for assessing visual fidelity, FVD for measuring temporal coherence, and CLIPSIM for detecting semantic alignment; (2) Motion Control Performance: Based on camera poses estimated by DUSt3R and object trajectories extracted by CoTrackerV3, motion control performance is quantified using CamMC and ObjMC metrics [57], which measure the Euclidean distance between predicted and ground-truth values; (3) Lighting Control Effectiveness: Evaluated through LPIPS, SSIM, and PSNR, comparing generated frames with ground-truth images to assess perceptual quality and structural fidelity.
4.2 Comparisons with State-of-the-Art Methods
Due to the lack of open-source I2V methods capable of simultaneously controlling camera and object motion, we evaluate VidCRAFT3 separately for camera motion control and object motion control, comparing it against SOTA methods in each domain.
[…]
Image Conductor 150.26 242.01 29.69 12.96
Motion-I2V 128.35 171.35 30.92 3.96
Ours 87.12 120.65 32.99 3.51
Table 2: Quantitative comparison with SOTA methods on WebVid-10M.
Camera Motion Control. VidCRAFT3 demonstrates superior performance in camera motion control compared to state-of-the-art (SOTA) methods like CameraCtrl, CamI2V, and MotionCtrl on the RealEstate 10K dataset. As shown in Table 1 and Fig. 5, VidCRAFT3 excels in camera motion control, as evidenced by its CamMC score of 4.07, which is lower than CameraCtrl (4.19), CamI2V (4.24), and MotionCtrl (4.23). VidCRAFT3 also achieves better results in other key metrics indicate that VidCRAFT3 not only controls camera motion more accurately but also generates videos with higher visual quality, temporal coherence, and semantic alignment. Qualitatively, VidCRAFT3 generates smoother and more realistic camera movements with fewer artifacts, particularly in complex scenes. This success stems from the Image2Cloud module, which enable precise 3D scene reconstruction and fine-grained camera control.
Object Motion Control. VidCRAFT3 demonstrates exceptional performance in object motion control, as highlighted in Table 2. The model achieves an ObjMC score of 3.51, which is lower than Image Conductor (12.96) and Motion-I2V (3.96). This indicates that VidCRAFT3 more accurately aligns the generated object trajectories with the ground truth, resulting in more realistic and faithful object motion. Additionally, VidCRAFT3 outperforms in other key metrics, showcasing superior visual quality, temporal coherence, and semantic alignment. Qualitatively, as shown in Fig. 6, VidCRAFT3 generates more realistic and consistent object movements compared to Image Conductor and Motion-I2V. The model effectively captures the dynamics of object motion, ensuring smooth transitions and natural interactions within the scene. These results highlight VidCRAFT3’s robust object motion control, driven by its advanced ObjMotionNet, which effectively captures and controls complex motion patterns.
[…]
CameraCtrl 3.4% 5.4% 4.7%
MotionCtrl 5.7% 7.3% 6.5%
CamI2V 5.6% 6.8% 5.2%
Ours 85.3% 80.5% 83.6%
[…]
Image Conductor 10.3% 11.9% 11.2%
Motion-I2V 14.8% 10.9% 13.2%
Ours 74.9% 77.2% 75.6%
Table 3: User study. Results demonstrating our method’s superior performance in camera and object motion control compared to baseline approaches across all metrics.
[…]
Text Cross-Atten 111.08 121.95 22.77 18.14 0.72 0.13 5.31
Time Embed. 101.71 123.31 22.70 19.07 0.73 0.12 5.21
Lighting Cross-Attn 100.83 117.69 23.70 19.49 0.74 0.11 5.00
Table 4: Ablation of lighting embedding integration strategies on VLD. Text Cross-Atten method concatenates lighting embedding with text embedding and integrates them into the model through text cross-attention. Time Embed method adds lighting embedding to time embedding. The proposed Lighting Cross-Atten method, detailed in Sec. 3.2, demonstrates enhanced performance in integrating lighting conditions.
User Study. To evaluate our method, we conducted a user study with 10 participants assessing 200 randomly selected results from our method and five baselines. The participants evaluated the results based on four key metrics: Camera Motion Precision, Object Motion Precision, Visual Quality, and Overall Quality. As shown in Table 3, our method achieved over 80% in camera motion control and over 74% in object motion control across all metrics, demonstrating precise and visually appealing control, validating its effectiveness and robustness in real-world applications.
4.3 Ablation Study
Training Strategy. We compare training with Dense, Sparse, and Dense+Sparse trajectories. As shown in Table 5, Dense training (ObjMC: 4.39) provides rich motion information but struggles during inference due to sparse inputs. Sparse training (ObjMC: 4.05) improves inference alignment but lacks motion detail. The Dense+Sparse approach achieves the best results (ObjMC: 3.51), leveraging dense trajectories for robust learning and fine-tuning with sparse trajectories for better generalization. This hybrid strategy enhances object motion control and video quality, as shown by superior FID (87.12) and FVD (120.65) scores.
Lighting Embedding Integration Strategies. We compare different lighting embedding integration strategies. As shown in Table 4, Lighting Cross-Atten achieves the best results, with an PSNR of 19.49, SSIM of 0.74, and LPIPS of 0.11, outperforming other methods. These findings confirm the effectiveness of Lighting Cross-Attention for precise lighting direction control.
Representation of Lighting Direction. We compare Fourier Embedding [31] and SH Encoding for representation of lighting direction. As shown in Table 6, SH Encoding outperforms Fourier Embedding, achieving better PSNR (19.49), SSIM (0.74), and LPIPS (0.11). These results confirm SH Encoding’s effectiveness for precise lighting control in image-to-video generation. For more results, please refer to the supplementary materials.
[…]
Dense 92.05 143.44 30.78 4.39
Sparse 91.54 123.15 30.93 4.05
Dense+Sparse 87.12 120.65 32.99 3.51
Table 5: Ablation of training strategy on WebVid-10M.
[…]
Fourier Embedding 107.40 121.89 21.71 17.48 0.70 0.14 5.03
SH Encoding 100.83 117.69 23.70 19.49 0.74 0.11 5.00
Table 6: Ablation of representation of lighting direction on VLD.
5 Conclusions
We present VidCRAFT3, a novel framework for precise image-to-video generation with simultaneous control over camera motion, object motion, and lighting direction. It integrates Image2Cloud for 3D point cloud generation, ObjMotionNet for encoding sparse object trajectories, and a Spatial Triple-Attention Transformer for incorporating lighting direction. The VideoLightingDirection dataset and a three-stage training strategy address real-world annotation challenges. Experiments show that VidCRAFT3 outperforms state-of-the-art methods in control granularity and visual coherence, advancing video generation toward more realistic and versatile applications.
Supplementary Material
Appendix A Preliminary
Image-to-video generation. Given a reference image $I_{\text{ref}}$ and some text description, image-to-video generation aims to generate a sequence of video frames $I=\{I_{1},I_{2},\ldots,I_{F}\}$ with $I_{1}=I$. In this work, we also allow the user to provide three forms of control signals: camera motion, object motion, and lighting direction. The video frames will be generated according to the text description and all the control signals provided. Specifically, the camera motion is represented by the movement of a set of camera extrinsics $E=\{E_{1},E_{2},\ldots,E_{F}\}$; the object motion is represented by a few sparse sequences of image pixels $\{s_{1}^{1},s_{1}^{2},s_{1}^{3},\ldots,s_{1}^{F}\},\{s_{1}^{2},s_{2}^{2},s_{2}^{3},\ldots,s_{2}^{F}\},\ldots$; the lighting direction is simply a unit 3d vector $L\in\boldsymbol{R}^{3}$.
Video Diffusion model (VDMs) represent a class of generative models that extend the principles of image diffusion to the domain of video generation. These models operate by defining a forward diffusion process that gradually transforms an initial video sample $x_{0}\sim p_{\text{data}}(x)$ into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ timesteps. The reverse process, parameterized by a denoising network $\epsilon_{\theta}(x_{t},t)$, learns to iteratively denoise the noisy latent representation $x_{t}$ to recover the original data $x_{0}$. The training objective is formulated as:
$\min_{\theta}\mathbb{E}_{t,x\sim p_{\text{data}},\epsilon\sim\mathcal{N}(0,I)}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2},$
where $\epsilon$ represents the ground truth noise, and $\theta$ denotes the learnable parameters of the network. Once trained, the model can generate high-quality videos by sampling from a random noise distribution $x_{T}$ and applying the learned denoising process iteratively.
To address the computational challenges associated with high-dimensional video data, Latent Diffusion Models (LDMs) are often employed. In this framework, a video $x\in\mathbb{R}^{F\times 3\times H\times W}$ is first encoded into a lower-dimensional latent space $\mathbf{z}=\mathcal{E}(x)$, where $\mathbf{z}\in\mathbb{R}^{F\times C\times h\times w}$. The diffusion and denoising processes are then performed in this latent space, significantly reducing computational complexity. The denoising process is conditioned on additional inputs $\mathbf{c}$, such as text prompts or motion control signals, enabling the generation of videos that adhere to specific semantic or temporal constraints. The final video is reconstructed through a decoder $\hat{x}=\mathcal{D}(\mathbf{z})$, ensuring both spatial and temporal coherence in the generated output.
Appendix B VLD Dataset Samples
Fig. 7 illustrates representative samples from our proposed VideoLightingDirection (VLD) Dataset, a synthetic dataset specifically designed to model complex light-object interactions. The dataset comprises two types of scenes: (a) Haven scenes, in which 3D models from Poly Haven are placed centrally within HDR environments; and (b) BOP scenes, featuring randomly positioned BOP models in six-plane textured rooms to simulate indirect lighting. Each subset includes video frames captured under two distinct lighting conditions while maintaining a consistent camera trajectory. This design highlights how changes in lighting direction impact object shading, reflections, and overall visual coherence, providing valuable data for training and evaluating models requiring precise lighting direction annotations.
Figure 7: Illustrations of samples from the proposed VideoLightingDirection (VLD) Dataset, featuring synthetic scenes designed to model complex light-object interactions. (a) Haven scenes, where 3D models from Poly Haven are placed at the center of HDR environments. (b) BOP scenes, with BOP models randomly positioned within six-plane textured rooms to simulate indirect lighting. Each subset includes video frames captured under two distinct lighting conditions, maintaining consistent camera trajectories.
Appendix C Qualitative Results of Ablation Study
In this section, we present qualitative results of the ablation studies described in the main text, providing visual insights into how different design choices impact the generated videos.
Training Strategy. Fig. 8 qualitatively compares videos generated using Dense, Sparse, and Dense+Sparse training strategies. Dense training results exhibit detailed object motion but struggle with precise alignment when tested with sparse trajectories. Sparse training achieves better alignment but at the cost of reduced motion detail. The Dense+Sparse approach effectively combines the strengths of both, delivering superior motion coherence and overall visual quality.
Lighting Embedding Integration Strategies. Fig. 9 compares the generated videos of different lighting embedding integration strategies. Lighting Cross-Attention demonstrates clear superiority, providing precise control over lighting directions, resulting in more realistic shadows, reflections, and overall lighting effects compared to Text Cross-Attention and Time Embedding.
Representation of Lighting Direction. Fig. 10 presents a visual comparison between Fourier Embedding and SH Encoding for representing lighting direction. SH Encoding consistently produces more realistic and detailed lighting effects, including enhanced shading and reflections, further supporting the quantitative findings in the main text.
These qualitative analyses further validate our key design decisions, emphasizing their impact on generating precise and realistic image-to-video translations under controlled camera motion, object motion, and lighting direction.
Appendix D Additional Qualitative Results
This section provides additional qualitative comparisons illustrating VidCRAFT3’s capabilities across multiple control dimensions.
Camera Motion + Lighting Direction Control. Fig. 11 demonstrates VidCRAFT3’s effectiveness in simultaneously controlling camera motion and lighting direction, showing clear advantages in lighting realism and trajectory accuracy compared to original sequences.
Camera Motion + Object Motion Control. Fig. 12 presents examples of simultaneous control over camera and object motions. VidCRAFT3 accurately reproduces intended object trajectories and maintains specified camera movements, closely matching the ground truth.
Camera Motion Control. Fig. 13 compares VidCRAFT3’s results with state-of-the-art methods (CameraCtrl, MotionCtrl, CamI2V) and ground-truth sequences, highlighting superior camera trajectory accuracy and visual coherence.
Object Motion Control. Fig. 14 provides comparisons against Image Conductor and Motion-I2V, showcasing VidCRAFT3’s superior capability in generating realistic object trajectories and visually coherent motion across diverse scenarios.
Lighting Direction Control. Fig. 15 further illustrates VidCRAFT3’s precise control over lighting direction, achieving accurate shading, shadows, and reflections, consistent with ground-truth sequences.
These additional qualitative evaluations highlight VidCRAFT3’s comprehensive and precise control capabilities, confirming its robustness and versatility across different control dimensions.
Appendix E Limitations
VidCRAFT3 can generate unstable results for many specific situations, e.g., large human motion, phyisical interactions, significant change of lighting conditions. This is mostly due to the lack of diverse training data and we believe the neural architecture can also be improved to promote a better understanding of physics and 3D spatial relationships of objects. The camera motion control and object motion control datasets contain inaccurate annotations of camera pose and motion trajectories, which sometimes result in blurred generated videos. Currently, VidCRAFT3 only offers control over the lighting direction and therefore can be improved to support full HDR or other more fine-grained representations of the light field.
Figure 8: Qualitative comparisons of training strategies (Dense, Sparse, Dense+Sparse) in the ablation study conducted on WebVid-10M. Results clearly illustrate that Dense training struggles with alignment given sparse input trajectories during inference, Sparse training improves alignment but loses motion detail, while Dense+Sparse training effectively combines both strategies, resulting in superior alignment and enhanced motion realism compared to the ground truth (GT).
Figure 9: Qualitative comparison of lighting embedding integration strategies (Text Cross-Attn, Time Embedding, Lighting Cross-Attn) in the ablation study on the VLD dataset. Results clearly demonstrate that the Lighting Cross-Attn strategy provides more accurate and realistic lighting control, achieving superior shadows, reflections, and overall visual fidelity compared to Text Cross-Attn and Time Embedding, aligning closely with the ground truth (GT).
Figure 10: Qualitative comparison between Fourier Embedding and SH Encoding for representing lighting direction on the VLD dataset. Results demonstrate that SH Encoding consistently produces more realistic lighting effects, achieving better shading accuracy and reflection details compared to Fourier Embedding, aligning closely with the ground truth (GT).
Figure 11: Additional qualitative results demonstrating simultaneous control over camera motion and lighting direction. Each pair of rows compares the original video with the corresponding relighting results generated by VidCRAFT3, highlighting the framework’s effectiveness in precisely adjusting lighting direction while preserving accurate camera motion.
Figure 12: Additional qualitative results showcasing simultaneous camera motion and object motion control. Each example compares our generated video (Ours) with ground-truth videos (GT), demonstrating VidCRAFT3’s capability to accurately reproduce intended object motion while precisely maintaining the specified camera motion.
Figure 13: Additional qualitative comparisons for camera motion control. Results from VidCRAFT3 (Ours) are compared with state-of-the-art methods (CameraCtrl, MotionCtrl, CamI2V) and ground-truth (GT) videos. These examples demonstrate VidCRAFT3’s superior capability in generating visually coherent and precise camera motion across diverse scenarios.
Figure 14: Additional qualitative comparisons for object motion control. Results generated by VidCRAFT3 (Ours) are compared against state-of-the-art methods (Image Conductor, Motion-I2V) and ground-truth (GT) videos. The examples demonstrate VidCRAFT3’s improved capability in accurately reproducing specified object trajectories, achieving more realistic and coherent object motion across diverse scenarios.
Figure 15: Additional qualitative results demonstrating lighting direction control. Our generated videos (”Ours”) are compared against ground-truth sequences (”GT”), clearly highlighting VidCRAFT3’s capability to accurately reproduce different lighting direction, achieving realistic shading, shadowing, and reflection details across diverse scenes.